A sea bird makes a sound as it dives along the surface of sea water at twenty degrees Celsius. A person hears the sound in room temperature air who is standing 15 meters away. A seal swimming directly below the sea bird also simultaneously hears the sound. What is the distance below the sea that the seal can be found?

a.

450m

b.

67m

c.

105m

d.

2.98m

e.

0.0436m
Answer #1

here,

at room temperature

the speed of sound in air , v1 = 343 m/s

the speed of sound in water , v2 = 1531 m/s

the distance of person , r1 = 15 m

let the distance of seal from the sea be v2

equating the time

r1 /v1 = r2 /v2

15 /343 = r2 /1531

r2 = 67 m

the distance of seal from the sea is b) 67 m